OpenStudy (anonymous):

find the second derivative of x^2 - y^2= 1

myininaya (myininaya):

2x-2yy'=0 solve for y' 2yy'=2x y'=2x/[2y]=x/y so y''=[(x)'(y)-(x)(y)']/y^2 y''=[1y-xy']/y^2 y''=[y-xy']/y^2 but y'=x/y y''=[y-x(x/y)]/y^2 y''=[y-x^2/y]/y^2 lets get rid of the double fraction y''=[y^2-x^2]/y^3 after multiply both top and bottom by y y''=-[x^2-y^2]/y^3 but omg x^2-y^2=1 so y''=-1/y^3 and done

OpenStudy (anonymous):

thank you so much! :)

myininaya (myininaya):

np